OpenStudy (loser66):

\(Let~~f:\mathbb R\rightarrow \mathbb R\) be a function. The graph of f, \(\large \Gamma =\{(x,y) \in \mathbb R^2: y=f(x)\}\). Suppose \(f\) is continous. 1) Prove that \(g:\mathbb R\rightarrow \mathbb R^2, g(x) =(x,f(x)) \)is continuous. 2)Prove that \(\Gamma_f\) is path connected. 3) Prove that \(\Gamma_f\) is closed in \(\mathbb R^2\) . Hint: prove that it contains all of its limit points. 4) Suppose \(\Gamma_f\) is closed and \(f\) is bounded. Prove that \(f\) is continuous. I need a big hand. Please, help.

OpenStudy (loser66):

I have class now. Please leave the guidance. I 'll be back later. Thanks in advance.

OpenStudy (anonymous):

I take it you know that f is continuous if and only if for any sequence \( x_n\) converging to\( x\) then\( f(x_n)\) converges to \( f(x)\)

OpenStudy (anonymous):

Another fact that we need is a bounded sequence \(z_n \) of reals has a convergent subsequence

OpenStudy (anonymous):

To prove that g is continuous, let \( x_n\) be a sequence converging to x, then \[ f(x_n) \to f(x) \implies g(x_n)= (x_n, f(x_n)) \to ( x,f(x))=g(x) \] Hence g is continuous

OpenStudy (anonymous):

2) The image of a path connected set by a continuous map is itself path connected.

OpenStudy (anonymous):

3) Let \[ (x_n, f(x_n)) \in \Gamma_f \quad \text { that converges to } (x,u) \] We need to show that u=f(x)

OpenStudy (anonymous):

Since f is continuous\[ f(x_n)\to f(x) \] but \[ f(x_n)\to u \] then \[ u=f(x) \]

OpenStudy (anonymous):

4) let \(x_n\) be a sequence converging to x, we need to prove that \( f(x_n) \to f(x) \). This can be done by proving that every \(x_{n_k} \)subsequence of of \(x_n\) has a further subsequence \(x_{n_{k_p}} \) such that \[ f(x_{n_{k_p}} )\to f(x) \] I would let you finish it now.

OpenStudy (loser66):

Thanks a lot. :)

OpenStudy (anonymous):

YW